(a) A company can manufacture a certain component from bar stock using an automatic lathe. The breakdown of costs is as follows: Cost of automatic lathe Tooling cost for component Setting up time Tool setters rate Machine operators rate Time to produce one component Cost of material per component General overheads (excluding $180,000.00 $ 2,000.00 10 hours $18.00 per hour $6.00 per hour 3 minutes $1.50 $15.00 per hour depreciation) Assume straight-line depreciation of the lathe over a period of 10 years, operating a 40 hour week for 50 weeks of the year. The anticipated life of the tooling is 1000 hours of production (i) If the company agrees to supply 18,500 components to a customer at a price of $3.50 per component, what profit, if any, will the company make on the job? (25 marks) (ii) What is the breakeven quantity for this job? (5 marks) (iii) Before production begins, the customer negotiates a revised contract. The new contract is for the supply of 22,500 components at a price per component of $3.20 What is the breakeven quantity for the new contract and what profit, if any, will the company make on the revised job? (25 marks) (b) Two projects are being assessed on their financial merit. The first project, Project A, involves an initial investment of $150,000. It is expected to generate an income of $30,000 per year for the next eight years. The income is generated at the end of each year. Project B does not require an initial lump sum investment but a series of annual payments. This project requires an annual fee of $35,000 to be paid at the start of each year for the first six years. Project B is then expected to generate a lump sum income of $340,000 after eight years have elapsed (45 marks) Assuming an interest rate of eight percent, calculate the net present value of each project and indicate which project represents the better financial investment. Present Value of $1 Year Discount Rate 8% 4 0.926 0681 0.735 Present Value of Annuity $1 4 0.794 0.540 Year Discount Rate 890 | 7 5.206 0.926 1.783 2.577 3.993 4.623 5.747
Answer #1
i. Per component 18500 components
Variable costs:
Cost of material 1.5 27750
Cost of Labour:
Tool setter(Per component--$ 18/60 min.*3 min.) 0.9 16650
Machine operator(Per component--$ 6/60 min. *3 min.) 0.3 5550
1.Total variable costs 2.7 49950
Fixed costs:
Tooling costs(2000/1000*925) 0.1 1850
General Overheads($ 15/hr.*925 hrs.) 0.75 13875
Depreciation-automatic lathe(180000/10yrs./2000 hrs.*925 hrs.) 0.45 8325
2.Total fixed costs 1.3 24050
Total costs for 18500 components(1+2) 4 74000
Offer price 3.5 64750
Loss 0.5 9250
NOTE: Share of full costs are considered as it is a Job order. Tool costs are indirect costs & hence treated as fixed costs.
ii.Breakeven quantity for the job:
Fixed costs/Contribution per component
ie.Fixed costs/(Offer price-Variable cost)per component
24050/(3.5-2.7)=
30063 components
iii.
Offer price/component 3.2
Less:Variable cost/component 2.7
Contribution /component 0.5
Total fixed costs (as in 1. above) 24050
Breakeven point=24050/(3.2-2.7)=
48100 components
b.
NPV (Project A):
-150000+(30000*5.747)=                       ( P/A 8%,8 yrs.)
22410
NPV (Project B):
(-35000*4.993)+(340000*0.540)=       (P/A due, 8%,6 yrs.)&P/F 8%,Yr.8)
8845
Project A represents better financial investment.